The piston diameter of a certain hand pump is 0.6 inch. The manager determines that the diameters are normally​ distributed, with a mean of 0.6 inch and a standard deviation of 0.004 inch. After recalibrating the production​ machine, the manager randomly selects 27 pistons and determines that the standard deviation is 0.0035 inch. Is there significant evidence for the manager to conclude that the standard deviation has decreased at the alpha equals 0.05 level of​ significance?
Answer #1

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: σ = 0.004
Alternative Hypothesis, Ha: σ < 0.004

Rejection Region
This is left tailed test, for α = 0.05 and df = 26
Critical value of Χ^2 is 15.379
Hence reject H0 if Χ^2 < 15.379

Test statistic,
Χ^2 = (n-1)*s^2/σ^2
Χ^2 = (27 - 1)*0.0035^2/0.004^2
Χ^2 = 19.906

P-value Approach
P-value = 0.204
As P-value >= 0.05, fail to reject null hypothesis.